if the current price of common stock is $55 per share in current dividends that was just paid was $2.20 per share, what is the required rate of return on the stock if the growth rate and dividend is expected to be 7% per year?
Answer #1

Current price = D0(1+g)/(rs-g)

55 = 2.20 (1+ .07)/(rs-.07)

55 = 2.20 * 1.07 /(rs-.07)

(rs- .07) = 2.20 *1.07 /55

(rs- .07) = 2.354/55

(rs- .07)   = .0428

rs =.0428 +.07

       = .1128 or 11.28%